2. A monopolist has a cost function given by TC -250+q+.004q2. The inverse market demand for boxes is given by p = governmental zoning rule. (a) What is its output and what price does it charge for boxes? (b) Calculate the firms profit at this output level. (c) Calculate the firms producers surplus at this output level. (d) Calculate the consumers surplus in this situation. 8-.001Q. The monopolist is currently able to exclude rivals from the market because of a special 3. The zoning rule is revoked, and the monopolist in problem 2 can no longer exclude others from using the same technology and producing boxes, so the market structure changes from monopoly to perfect competition. (That is, assume that all firms are price-takers and that they produce at minimum average cost in equilibrium.) (a) What will the market price and quantity be? (b) Calculate consumers surplus under this market structure. (c) Calculate aggregate producer profits and producers surplus. (d) Comparing the two market structures for this industry, which has a larger social surplus (sum of producers and consumers surplus)?
Answer #1

2)

Total Cost (TC) = 250+ q +0.004q2

Demand: p = 8 - 0.001Q

a) The monopolist will produce where the marginal revenue equals the marginal cost.

MC = dTC/dq

MC = 1+0.008q

TR = P*Q

TR = 8Q – 0.001Q2

Marginal Revenue(MR) = dTR/dQ

MR = 8-0.002Q

Therefore,

1+0.008q = 8 – 0.002q

0.01q = 7

q = 700

Price = 8 – 0.001*700

Price = 7.3

b) Profit = TR – TC

Profit = 7.3*700 – (250+ 700+1960)

Profit = 5110 - 2910 = $2200

c) Producer surplus = Profit = $2200

d) Consumer Surplus = 0.5*(8 – 7.3)*700 = $245

*We are supposed to do only one question.